A 12.2 g sample of an aqueous solution of perchloric acid contains an unknown amount of the acid.
If 13.7 mL of 0.419 M barium hydroxide are required to neutralize the perchloric acid, what is the percent by mass of perchloric acidin the mixture?

_____ % by mass
Answer #1

The balanced reaction will be

Ba(OH)2(aq) + 2HClO4(aq) ---------------------- Ba(ClO4)2 + H2O

Number of moles of Barium Hydroxide = Volume of solution (in L) * Molarity(M) = 13.7/1000 * 0.419 = 0.00574 moles

1 mole of Ba(OH)2 reacts with 2 moles of HClO4

Hence, moles of HClO4 will be 2 * 0.00574 = 0.01148 moles

Molar mass of HClO4 = 100.46 g/mol

Mass of HClO4 in the mixture = Number of moles * Molar mass = 100.46 g/mol * 0.01148 mol = 1.153 grams

Percent by mass in mixture = 1.153/12/2 * 100 = 4.81%